[Brahms-l] Shift report 20050425 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Apr 25 2005 - 23:50:48 EDT
Shifter leader was kh
Shifters were sk


We had beam from ~21:00 and the store is still in progress estimated to dump at 03:00.
Runs achieved.
Run  tr2   tr3
FS 8 deg 3450A
14274 238  130k
14275 111  56k
The absence of k on the trigger 2 counts is not an accident.  Decide that the count rate is too low, so we change the field to 1/sqrt(2)
FS 8 deg 2442A
14276 in progress


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Apr 25 23:51:04 2005

This archive was generated by hypermail 2.1.8 : Mon Apr 25 2005 - 23:51:21 EDT